OpenStudy (anonymous):

using substitution method need help solving this problem step by step please. y=x+3 y=5x-5

OpenStudy (anonymous):

substitute y=x+3 in the 2nd equation=> x+3=5x-5=>3+5=5x-x=>8=4x=>x=2 y=x+3=2+3=5